# Revisiting Isocurvature Bounds on the Minimal QCD Axion

### v1

@@ 4.2 Numerical solution across the parameter space
    In summary, a detailed reconstruction of the dynamics and decay of the saxion energy density (as e.g. in [93, 39]) can clarify the cosmological evolution and, up to some model dependence, might identify some regions of parameter space that are not allowed. We find that the issue of saxion decay does not pose a problem over most of the points of our scan, and does not generically challenge the possibility to avoid isocurvature via an enhancement of $f_{a}^{\text{(inf)}}$.
@@ 5 Conclusions
The isocurvature bounds are highly UV dependent. While it is possible for the region shown in Fig. 2 to overproduce isocurvature perturbations, it does not have to happen. Whether or not a particular point is ruled out depends on UV parameters of the PQ-breaking potential as well as the models of inflation and reheating. The minimal model for the PQ-breaking potential has the form shown in Eq. 3.2. When $\lambda_{\Phi}$ is small and $\xi\sim\mathcal{O}(1)$ (or really the ratio $\lambda_{\Phi}/\xi\ll 1$) then during inflation the axion will have a much larger effective PQ-breaking scale $f_{a}^{\text{(inf)}}$ than it will in the later universe. This automatically suppresses the isocurvature perturbations.
While this was realized previously, it was believed that this possibility was ruled out because the large drop in the saxion field $\phi$ at the end of inflation would drive parametric resonance of the axion, creating axion strings and thus overproducing axion dark matter over most of this region. We have shown that in fact this does not necessarily happen. There are many points in the UV parameter space of this model where parametric resonance is weak and the axion is not overproduced. The strength of this parametric resonance depends both on the UV parameters of the minimal QCD axion model in Eq. 3.2 as well as the model of inflation and reheating. We chose two regular inflation models and demonstrated that many parts of the parameter space work in each model. The exact region in the $f_{a}$-$H_{I}$ plane which is rescued from isocurvature this way depends on the models of inflation and rehehating as shown in Fig. 6.
Figure 7: Summary plot of the region (shaded in grey) that is not rescued from the isocurvature constraint even by an enhancement of $f_{a}^{\text{(inf)}}$ during inflation in the inflation models we chose. This plot assumes a threshold $f_{a}^{\text{(inf)}}/f_{a}<R_{f_{a}}$ (with $R_{f_{a}}=700$) for which parametric resonance of the axion can be avoided, as given in Eq. 4.12. Note that a different inflation model or a different UV model of the axion might cause even the grey shaded regions to be allowed and thus we cannot say conclusively that they are ruled out by isocurvature. The reheating line (in orange) assumes $T_{\text{max}}=T_{\mathrm{inst.\,RH}}/\sqrt{10}$.
@@
Acknowledgments
We thank Marco Gorghetto, Ed Hardy, Liam McAllister, Riccardo Natale, Maximilian Ruhdorfer, Ben Safdi, and Géraldine Servant for helpful discussions.
The authors acknowledge support by NSF Grant PHY-2310429, Simons Investigator Award No. 824870, DOE HEP QuantISED award #100495, the Gordon and Betty Moore Foundation Grant GBMF7946. D.R. is supported at U. of Zurich by the UZH Postdoc Grant 2023 Nr. FK-23-130.
@@ A.1 Examples of parametric resonance for the axion
This model-dependent part of the dynamics determines the position and velocity of the saxion background (green line) as visible in the top left corner of Figs. 8 and 9. The blue lines mark the positions of the saxion minima (in the direction of the initial misalignment angle $\theta_{\text{i}}$) as a function of time. For the warm inflation model considered in this section, at the end of inflation they relax exponentially fast to their late-time value $f_{a}$.
Figure 8: Evolution of a point in parameter space displaying a strong parametric resonance. As the saxion (shown in green) oscillates widely around the origin, the axion (shown in purple) direction grows exponentially, until the saxion settles in a minimum and its amplitude decreases. Top left: time evolution of the saxion (radial direction; strictly speaking, this panel shows $\overline{X}(t)$) after $t_{e}$ (marked by a vertical red line; time is in units of $H_{e}^{-1}$) in green, with the blue lines showing the position of the minima of the potential. Right: Spectrum of the growth of axion perturbations due to parametric resonance, computed as the maximum amplitude reached by each mode $y_{k}\approx\delta a_{k}(t)$ normalised to its initial value at $t_{e}$. Bottom left: variance of axion fluctuations (more precisely, we plot the fluctuations $\langle(Y-\overline{Y})^{2}\rangle$, which match the axion power spectrum up to a factor $(\phi(t)/f_{a})^{2}$), normalised to $(\pi f_{a})^{2}$. The dashed red line marks the threshold for the formation of axion strings, and the restoration of the PQ symmetry. When the axion passes this threshold, our linearised analysis is no longer a good approximation.
@@
In the right-hand side panels of Figs. 8 and 9, we plot the maximum enhancement reached by $\sim 250$ modes $y_{k}$ in the range $(0.2-10)k_{e}$, with $k_{e}=a_{e}H_{e}$, evolved according to the linearised equations of motion Eq. 4.7b in the background evolution of $\overline{X}(t)$. The initial amplitude of $y_{k}$ around Hubble crossing is set by quantum fluctuations, as discussed in Section 4.2. As these modes pass through the first instability band of the parametric resonance, they grow exponentially for some time. The unstable modes, in units of $k_{e}$, are given by $\phi(t)\cdot\sqrt{\lambda_{\Phi}}/H(t)$, which can be read off by the vertical range of the top left panels.
The power spectrum $\tfrac{1}{\pi^{2}f_{a}^{2}}\int\tfrac{k^{3}}{2\pi^{2}}|y_{k}|^{2}\mathrm{d}\ln k=\tfrac{1}{\pi^{2}}\langle(\theta(x)-\theta_{\text{i}})^{2}\rangle$ is plotted in the bottom left panels of Figs. 8 and 9. As this quantity becomes $\gtrsim\mathcal{O}(10^{-2})$, our linearised analysis ceases being a good approximation, as the terms that we neglected in deriving Eq. 4.7 become large, and the cross-coupling of various modes is non negligible, requiring a lattice study. For our scope, a linearised analysis is perfectly suitable to identify a safe range where the non-thermal symmetry restoration is avoided and strings are not formed. We notice that, for a given $k$, the oscillations of $\delta y_{k}^{2}$ do not depend on our choice of the phase of its initial conditions (as we checked explicitly), but they are in phase with the radial mode which is driving the resonance. By integrating over various modes in the power spectrum, these synchronous oscillations of different modes average out only partially, as visible in the bottom left panels of Figs. 8 and 9, and they are present as long as the parametric resonance is active.
We notice, from the comparison of the top and bottom left panels of Figs. 8 and 9, that the parametric resonance occurs for about as long as the saxion field probes the concave part of its potential $V(\phi)$ (shaded in light red in the figures), so that the frequency of some axion fluctuations $y_{k}$ is negative in Eq. 4.7b.
@@
$\delta\theta_{\text{(par.res.)}}\sim\frac{H_{I}}{2\pi f_{a}}\exp\left(\mathcal{O}(1)\cdot n_{\text{$\phi$-crossings}}\right)\,,$ (A.1)
with $n_{\text{$\phi$-crossings}}$ defined as the number of times that the saxion crossed the origin (respectively 21 and 8 in Figs. 8 and 9). This number depends in turn on $f_{a}^{\text{(inf)}}/f_{a}$, as the evolution of $f_{a}^{\text{(inf)}}(t)$ towards the end of inflation determines the initial amplitude of the saxion when it is released above the wine-bottle potential. This relation is more difficult to estimate analytically though, as it depends on the detailed expression of $H_{I}(t)$ towards the end of inflation. $n_{\text{$\phi$-crossings}}$ might also be enhanced from oscillations of $R(t)$ in the preheating phase, before the universe is radiation-dominated and $R\approx 0$.
After having illustrated in detail the result of our numerical analysis for two points of the scan, we plot in Fig. 10 the points of our scan for the model of Warm Inflation projected on the 2-dimensional subspaces of the variables $(f_{a},\lambda_{\Phi},\xi)$ (while Fig. 3 shows them in the $(H_{I},f_{a})$ plane). Analogously to that figure, the colour code of the points marks in blue the points where $\delta\theta_{\text{(par.res.)}}\lesssim\mathcal{O}(10^{-3})$ and strings are not formed.
@@ A.2 Sources of friction for saxion and axion
In this appendix, we only comment on two possible scenarios of such friction sources. A first channel, that is guaranteed for the QCD axion, is the sphaleron friction into SM gluons for the axion [94, 95, 96, 97]. Accounting for the presence of a chiral suppression of the sphaleron emission rate in presence of light quarks [94, 98, 99] we find that, also due to the decrease of $\alpha_{\text{s}}$ at high scales around $T\sim 10^{13}\,\text{GeV}$, this friction is of order $\Gamma_{\text{sph}}\approx\tfrac{(N_{c}\alpha_{\text{s}})^{5}T^{3}}{2f_{a}^{2}}\frac{y_{q}^{2}}{2(N_{c}\alpha_{\text{s}})^{4}+y_{q}^{2}}\approx 10^{-14}T\big{(}\tfrac{y_{q}}{10^{-4}}\tfrac{T/f_{a}}{10^{-2}}\big{)}^{2}\ll H$. This friction channel is then negligible in a minimal scenario.
@@
Another scenario for which we perform some estimates, summarised in Fig. 11, is a KSVZ completion for the axion, with coloured fermions $\psi$ coupled to the saxion field via a Yukawa coupling $y_{\textsc{ksvz}}$. We also understand the common assumption that these KSVZ fermions have a hypercharge that allows them to decay into SM quarks, to avoid the risk of overclosure. The first orange band shows where thermal corrections from a bath of KSVZ fermions do not restore $U(1)_{\textsc{pq}}$: their thermal corrections might be subdominant (first range), or too heavy to be produced and stable for enough time to effectively restore the symmetry (second range). The second pink band quantifies where the 1-loop quantum corrections $\delta\lambda_{\Phi}\sim\tfrac{N_{c}}{(4\pi)^{2}}y_{\textsc{ksvz}}^{4}$ are smaller than $\lambda_{\Phi}$, for the points considered in our scan. The third blue band quantifies the rough range where we found that the decay $\phi\to\overline{\psi}\psi$ (kinematically open when $y_{\textsc{ksvz}}f_{a}<T$, see the detailed discussion in [93, 39]) could impact appreciably the evolution of the saxion, reducing the growth of the axion modes. In summary, from our rough estimates there appears to be window of opportunity for $y_{\textsc{ksvz}}\sim 10^{-3}-10^{-2}$ where the KSVZ fermions might induce a non-negligible friction for the saxion and reduce the parametric resonance of the axion, while not restoring $U(1)_{\textsc{pq}}$.
@@ Appendix B Comparison of different inflationary models and couplings to the saxion
Figure 12: Evolution of the quantities $(R,\dot{\sigma}^{2},\sigma^{2})$ ($\sigma$ being the inflaton field) that can induce a negative mass term for the saxion during inflation (enhancing $f_{a}^{\text{(inf)}}$) through couplings $\xi R|\Phi|^{2}$, $(\partial_{\mu}\sigma)^{2}|\Phi|^{2}$, $\sigma^{2}|\Phi|^{2}$. The quantities are normalised to their initial value. The green and red vertical lines in the four plots mark respectively the times at $60$ $e$-folds before $t_{e}$, and $t_{e}$. The four panels corresponds to four inflationary models: a Warm inflation model, Starobinsky inflation, quadratic and hilltop inflation. The models on the top row are the ones for which we performed a scan of the saxion parameter space.
The model for which we have displayed most of our results is a Warm Inflation model, as it displays a wider decrease of $H_{I}(t)$ during the 60 $e$-folds of inflation and the end of inflation displays a smooth transition to radiation domination with $R=0$, without intermediate matter dominated phase where $R$ oscillates (adding a further layer of complexity to the evolution of saxion and axion). As a concrete implementation, we assume that the inflaton $\sigma$ couples with a strongly coupled Yang-Mills dark sector via a coupling $\sigma G\widetilde{G}$, inducing a friction for the inflaton due to thermal production of sphalerons [94, 95, 96, 97]. The effective friction term in the inflaton equations of motion becomes $(3H+\Upsilon_{\text{sph}})\dot{\sigma}$, with
$\Upsilon_{\text{sph}}$ $=(N_{c}\alpha)^{5}\frac{T_{\textsc{dr}}^{3}}{2f^{2}}\,,$ $N_{c}=3,\ \alpha=0.1,$ (B.1)
@@
with $f=1.17\cdot 10^{13}\,\text{GeV}\cdot(H_{I}/10^{10}\,\text{GeV})^{1/4}$ chosen to ensure a regime of strong warm inflation with $Q\equiv\Upsilon_{\text{sph}}/(3H_{I})\approx 180$. We notice that this value lies within 1-2 orders of magnitude from the axion decay constants $f_{a}$ of most points of our scan in Fig. 3, which can ease the formulation of a complete model for saxion and inflaton. As for the inflationary potential, we use the Starobinsky potential in Eq. B.2 with $c=0.3$, which was studied in a warm inflation setting in [96]. Finally, we notice that the bath at a temperature $T_{\textsc{dr}}$ might not include the SM species if these are not appreciably coupled to the dark Yang-Mills sector. The reheat temperature of the SM and PQ sectors determine the boundary of the post-inflationary region in Figs. 3 and 5, and it might be significantly smaller than $T_{\textsc{dr}}(t_{e})$. As a conservative assumption, we take them to coincide, but it is worth remembering that the lower boundary of the “rescued” region in Fig. 5 might be lowered appreciably.
The second model for which we perform a scan of the parameter space is Starobinsky inflation [100] ($R+R^{2}$), which is also a special case $\alpha=1$ of the $\alpha$-attractor models [101, 102, 103]. We use the scalar-field version of the Starobinsky model (related via a conformal transformation [104]), and the inflaton potential reads
$V_{\text{Starobinsky}}(\sigma)=3H_{I}^{2}M_{\textsc{p}}^{2}\left(1-\exp\left(-c\frac{\sigma}{M_{\textsc{p}}}\right)\right)^{2}\,,\quad c=\sqrt{2/3}\,.$ (B.2)
We model the reheating phase with a decay rate $\Gamma_{\sigma}=10^{-(1,3,5)}m_{\sigma}$, with $m_{\sigma}=2H_{I}$ the inflaton mass in the origin. Starting from the time $t_{e}$, we turn this decay rate on in the equations of motion for the inflaton, with a friction $(3H+\Gamma_{\sigma})\dot{\sigma}$. By comparing the evolution of $R$ in Fig. 12 to the Warm Inflation case, this Starobinsky model displays a shorter and faster excursion of $H_{I}(t)$ in the last few $e$-folds, which contributes to making the evolution of the saxion less adiabatic and the axion field more subject to parametric resonance (see Fig. 6). This point, together with the persistence of an oscillating $R(t)$ in the preheating phase, reduce the ratio $R_{f_{a}}=f_{a}^{\text{(inf)}}/f_{a}$ to avoid string formation (Eq. 4.12) from the value of 700 for Warm Inflation to roughly $(60,120,180)$ for $\Gamma_{\sigma}=10^{-(1,3,5)}m_{\sigma}$. The analogue plots of Fig. 5 for the Starobinsky model are reproduced in Fig. 13.

### v2

@@ 4.2 Numerical solution across the parameter space
    In summary, a detailed reconstruction of the dynamics and decay of the saxion energy density (as e.g. in Moroi:2013tea; Moroi:2014mqa; Gouttenoire:2021jhk; Eroncel:2024rpe) can clarify the cosmological evolution and, up to some model dependence, might identify some regions of parameter space that are not allowed. We find that the issue of saxion decay does not pose a problem over most of the points of our scan, and does not generically challenge the possibility to avoid isocurvature via an enhancement of $f_{a}^{\text{(inf)}}$.
@@ 5 Conclusions
The isocurvature bounds are highly UV dependent. While it is possible for the region shown in Fig. 2 to overproduce isocurvature perturbations, it does not have to happen. Whether or not a particular point is ruled out depends on UV parameters of the PQ-breaking potential as well as the models of inflation and reheating. The minimal model for the PQ-breaking potential has the form shown in Eq. 10. When $\lambda_{\Phi}$ is small and $\xi\sim\mathcal{O}(1)$ (or really the ratio $\lambda_{\Phi}/\xi\ll 1$) then during inflation the axion will have a much larger effective PQ-breaking scale $f_{a}^{\text{(inf)}}$ than it will in the later universe. This automatically suppresses the isocurvature perturbations.
While this was realized previously, it was believed that this possibility was ruled out because the large drop in the saxion field $\phi$ at the end of inflation would drive parametric resonance of the axion, creating axion strings and thus overproducing axion dark matter over most of this region. We have shown that in fact this does not necessarily happen. There are many points in the UV parameter space of this model where parametric resonance is weak and the axion is not overproduced. The strength of this parametric resonance depends both on the UV parameters of the minimal QCD axion model in Eq. 10 as well as the model of inflation and reheating. We chose two regular inflation models and demonstrated that many parts of the parameter space work in each model. The exact region in the $f_{a}$-$H_{I}$ plane which is rescued from isocurvature this way depends on the models of inflation and rehehating as shown in Fig. 6.
Figure 8: Summary plot of the region (shaded in grey) that is not rescued from the isocurvature constraint even by an enhancement of $f_{a}^{\text{(inf)}}$ during inflation in the inflation models we chose. This plot assumes a threshold $f_{a}^{\text{(inf)}}/f_{a}<R_{f_{a}}$ (with $R_{f_{a}}=700$) for which parametric resonance of the axion can be avoided, as given in Eq. 24. Note that a different inflation model or a different UV model of the axion might cause even the grey shaded regions to be allowed and thus we cannot say conclusively that they are ruled out by isocurvature. The reheating line (in orange) assumes $T_{\text{max}}=T_{\mathrm{inst.\,RH}}/\sqrt{10}$.
@@
Acknowledgements.
We thank Marco Gorghetto, Ed Hardy, Junwu Huang, Liam McAllister, Riccardo Natale, Maximilian Ruhdorfer, Ben Safdi, Géraldine Servant and Sebastian Zell for helpful discussions. The authors acknowledge support by NSF Grant PHY-2310429, Simons Investigator Award No. 824870, DOE HEP QuantISED award #100495, the Gordon and Betty Moore Foundation Grant GBMF7946. D.R. is supported at U. of Zurich by the UZH Postdoc Grant 2023 Nr. FK-23-130.
@@ A.1 Examples of parametric resonance for the axion
This model-dependent part of the dynamics determines the position and velocity of the saxion background (green line) as visible in the top left corner of Figs. 9 and 10. The blue lines mark the positions of the saxion minima (in the direction of the initial misalignment angle $\theta_{\text{i}}$) as a function of time. For the warm inflation model considered in this section, at the end of inflation they relax exponentially fast to their late-time value $f_{a}$.
Figure 9: Evolution of a point in parameter space displaying a strong parametric resonance. As the saxion (shown in green) oscillates widely around the origin, the axion (shown in purple) direction grows exponentially, until the saxion settles in a minimum and its amplitude decreases. Top left: time evolution of the saxion (radial direction; strictly speaking, this panel shows $\overline{X}(t)$) after $t_{e}$ (marked by a vertical red line; time is in units of $H_{e}^{-1}$) in green, with the blue lines showing the position of the minima of the potential. Right: Spectrum of axion perturbations due to parametric resonance, computed as the maximum amplitude reached by each mode $y_{k}\approx\delta a_{k}(t)$. Bottom left: variance of axion fluctuations (more precisely, we plot the fluctuations $\langle(Y-\overline{Y})^{2}\rangle$, which match the axion power spectrum up to a factor $(\phi(t)/f_{a})^{2}$), normalised to $(\pi f_{a})^{2}$. The dashed red line marks the threshold for the formation of axion strings, and the restoration of the PQ symmetry. When the axion passes this threshold, our linearised analysis is no longer a good approximation.
@@
In the right-hand side panels of Figs. 9 and 10, we plot the maximum enhancement reached by $\sim 250$ modes $y_{k}$ in the range $(0.2-10)k_{e}$, with $k_{e}=a_{e}H_{e}$, evolved according to the linearised equations of motion Eq. 19b in the background evolution of $\overline{X}(t)$. We sample three random initial phases for $y_{k}$ and we average over the corresponding solutions. The initial amplitude of $y_{k}$ around Hubble crossing is set by quantum fluctuations, as discussed in Section 4.2. As these modes pass through the first instability band of the parametric resonance, they grow exponentially for some time. The unstable modes, in units of $k_{e}$, are given by $\phi(t)\cdot\sqrt{\lambda_{\Phi}}/H(t)$, which can be read off by the vertical range of the top left panels.
The power spectrum $\tfrac{1}{\pi^{2}f_{a}^{2}}\int\tfrac{k^{3}}{2\pi^{2}}|y_{k}|^{2}\mathrm{d}\ln k=\tfrac{1}{\pi^{2}}\langle(\theta(x)-\theta_{\text{i}})^{2}\rangle$ is plotted in the bottom left panels of Figs. 9 and 10. As this quantity becomes $\gtrsim\mathcal{O}(10^{-2})$, our linearised analysis ceases being a good approximation, as the terms that we neglected in deriving Eq. 19 become large, and the cross-coupling of various modes is non negligible, requiring a lattice study. For our scope, a linearised analysis is perfectly suitable to identify a safe range where the non-thermal symmetry restoration is avoided and strings are not formed. We notice that, for a given $k$, the oscillations of $\delta y_{k}^{2}$ do not depend on our choice of the phase of its initial conditions (as we checked explicitly), but they are in phase with the radial mode which is driving the resonance. By integrating over various modes in the power spectrum, these synchronous oscillations of different modes average out only partially, as visible in the bottom left panels of Figs. 9 and 10, and they are present as long as the parametric resonance is active.
We notice, from the comparison of the top and bottom left panels of Figs. 9 and 10, that the parametric resonance occurs for about as long as the saxion field probes the concave part of its potential $V(\phi)$ (shaded in light red in the figures), so that the frequency of some axion fluctuations $y_{k}$ is negative in Eq. 19b.
@@
$\delta\theta_{\text{(par.res.)}}\sim\frac{H_{I}}{2\pi f_{a}}\exp\left(\mathcal{O}(1)\cdot n_{\text{$\phi$-crossings}}\right)\,$, (26)
with $n_{\text{$\phi$-crossings}}$ defined as the number of times that the saxion crossed the origin (respectively 21 and 8 in Figs. 9 and 10). This number depends in turn on $f_{a}^{\text{(inf)}}/f_{a}$, as the evolution of $f_{a}^{\text{(inf)}}(t)$ towards the end of inflation determines the initial amplitude of the saxion when it is released above the wine-bottle potential. This relation is more difficult to estimate analytically though, as it depends on the detailed expression of $H_{I}(t)$ towards the end of inflation. $n_{\text{$\phi$-crossings}}$ might also be enhanced from oscillations of $R(t)$ in the preheating phase, before the universe is radiation-dominated and $R\approx 0$.
@@ A.2 Sources of friction for saxion and axion
In this appendix, we only comment on two possible scenarios of such friction sources. A first channel, that is guaranteed for the QCD axion, is the sphaleron friction into SM gluons for the axion McLerran:1990de; Berghaus:2019whh; DeRocco:2021rzv; Mirbabayi:2022cbt. Accounting for the presence of a chiral suppression of the sphaleron emission rate in presence of light quarks McLerran:1990de; Berghaus:2020ekh; Berghaus:2025dqi we find that, also due to the decrease of $\alpha_{\text{s}}$ at high scales around $T\sim 10^{13}\,\text{GeV}$, this friction is of order $\Gamma_{\text{sph}}\approx\tfrac{(N_{c}\alpha_{\text{s}})^{5}T^{3}}{2f_{a}^{2}}\frac{y_{q}^{2}}{2(N_{c}\alpha_{\text{s}})^{4}+y_{q}^{2}}\approx 10^{-14}T\big(\tfrac{y_{q}}{10^{-4}}\tfrac{T/f_{a}}{10^{-2}}\big)^{2}\ll H$. This friction channel is then negligible in a minimal scenario.
@@
Another scenario for which we perform some estimates, summarised in Fig. 11, is a KSVZ completion for the axion, with coloured fermions $\psi$ coupled to the saxion field via a Yukawa coupling $y_{\textsc{ksvz}}$. We also understand the common assumption that these KSVZ fermions have a hypercharge that allows them to decay into SM quarks, to avoid the risk of overclosure. The first orange band shows where thermal corrections from a bath of KSVZ fermions do not restore $U(1)_{\textsc{pq}}$: their thermal corrections might be subdominant (first range), or too heavy to be produced and stable for enough time to effectively restore the symmetry (second range). The second pink band quantifies where the 1-loop quantum corrections $\delta\lambda_{\Phi}\sim\tfrac{N_{c}}{(4\pi)^{2}}y_{\textsc{ksvz}}^{4}$ are smaller than $\lambda_{\Phi}$, for the points considered in our scan. The third blue band quantifies the rough range where we found that the decay $\phi\to\overline{\psi}\psi$ (kinematically open when $y_{\textsc{ksvz}}f_{a}<T$, see the detailed discussion in Gouttenoire:2021jhk; Eroncel:2024rpe) could impact appreciably the evolution of the saxion, reducing the growth of the axion modes. In summary, from our rough estimates there appears to be window of opportunity for $y_{\textsc{ksvz}}\sim 10^{-3}-10^{-2}$ where the KSVZ fermions might induce a non-negligible friction for the saxion and reduce the parametric resonance of the axion, while not restoring $U(1)_{\textsc{pq}}$.
@@ Appendix B Comparison of different inflationary models and couplings to the saxion
Figure 12: Evolution of the quantities $(R,\dot{\sigma}^{2},\sigma^{2})$ ($\sigma$ being the inflaton field) that can induce a negative mass term for the saxion during inflation (enhancing $f_{a}^{\text{(inf)}}$) through couplings $\xi R|\Phi|^{2}$, $(\partial_{\mu}\sigma)^{2}|\Phi|^{2}$, $\sigma^{2}|\Phi|^{2}$. The quantities are normalised to their initial value. The green and red vertical lines in the four plots mark respectively the times at 6060 $e$-folds before $t_{e}$, and $t_{e}$. The four panels corresponds to four inflationary models: a Warm inflation model, Starobinsky inflation, quadratic and hilltop inflation. The models on the top row are the ones for which we performed a scan of the saxion parameter space.
The model for which we have displayed most of our results is a Warm Inflation model, as it displays a wider decrease of $H_{I}(t)$ during the 60 $e$-folds of inflation and the end of inflation displays a smooth transition to radiation domination with $R=0$, without intermediate matter dominated phase where $R$ oscillates (adding a further layer of complexity to the evolution of saxion and axion). As a concrete implementation, we assume that the inflaton $\sigma$ couples with a strongly coupled Yang-Mills dark sector via a coupling $\sigma G\widetilde{G}$, inducing a friction for the inflaton due to thermal production of sphalerons McLerran:1990de; Berghaus:2019whh; DeRocco:2021rzv; Mirbabayi:2022cbt. The effective friction term in the inflaton equations of motion becomes $(3H+\Upsilon_{\text{sph}})\dot{\sigma}$, with
[…]
@@
with $f=1.17\cdot 10^{13}\,\text{GeV}\cdot(H_{I}/10^{10}\,\text{GeV})^{1/4}$ chosen to ensure a regime of strong warm inflation with $Q\equiv\Upsilon_{\text{sph}}/(3H_{I})\approx 180$. We notice that this value lies within 1-2 orders of magnitude from the axion decay constants $f_{a}$ of most points of our scan in Fig. 3, which can ease the formulation of a complete model for saxion and inflaton. As for the inflationary potential, we use the Starobinsky potential in Eq. 28 with $c=0.3$, which was studied in a warm inflation setting in DeRocco:2021rzv. Finally, we notice that the bath at a temperature $T_{\textsc{dr}}$ might not include the SM species if these are not appreciably coupled to the dark Yang-Mills sector. The reheat temperature of the SM and PQ sectors determine the boundary of the post-inflationary region in Figs. 3 and 5, and it might be significantly smaller than $T_{\textsc{dr}}(t_{e})$. As a conservative assumption, we take them to coincide, but it is worth remembering that the lower boundary of the “rescued” region in Fig. 5 might be lowered appreciably.
The second model for which we perform a scan of the parameter space is Starobinsky inflation Starobinsky:1980te ($R+R^{2}$), which is also a special case $\alpha=1$ of the $\alpha$-attractor models Ferrara:2013rsa; Kallosh:2013tua; Kehagias:2013mya. We use the scalar-field version of the Starobinsky model (related via a conformal transformation Whitt:1984pd), and the inflaton potential reads
$V_{\text{Starobinsky}}(\sigma)=3H_{I}^{2}M_{\textsc{p}}^{2}\left(1-\exp\left(-c\frac{\sigma}{M_{\textsc{p}}}\right)\right)^{2}\,,\quad c=\sqrt{2/3}\,.$ (28)
We model the reheating phase with a decay rate $\Gamma_{\sigma}=10^{-(1,3,5)}m_{\sigma}$, with $m_{\sigma}=2H_{I}$ the inflaton mass in the origin. Starting from the time $t_{e}$, we turn this decay rate on in the equations of motion for the inflaton, with a friction $(3H+\Gamma_{\sigma})\dot{\sigma}$. By comparing the evolution of $R$ in Fig. 12 to the Warm Inflation case, this Starobinsky model displays a shorter and faster excursion of $H_{I}(t)$ in the last few $e$-folds, which contributes to making the evolution of the saxion less adiabatic and the axion field more subject to parametric resonance (see Fig. 6). This point, together with the persistence of an oscillating $R(t)$ in the preheating phase, reduce the ratio $R_{f_{a}}=f_{a}^{\text{(inf)}}/f_{a}$ to avoid string formation (Eq. 24) from the value of 700 for Warm Inflation to roughly $(60,120,180)$ for $\Gamma_{\sigma}=10^{-(1,3,5)}m_{\sigma}$. The analogue plots of Fig. 5 for the Starobinsky model are reproduced in Fig. 13.